The prevalence of undetected diabetes in a population to be screened is approximately 1.5%, and it is assumed that 10,000 persons will be screened. The screening test will measure blood serum glucose content. A value of 180 mg% or higher is considered positive. The sensitivity and specificity of associated with this screening test are 22.9% and 99.8%, respectively . 10. Construct a 2x2 table.

What is the positive predictive value?

What is the negative predictive value?
Answer #1

below is 2*2 table

tested positive tested negative total
have disease 150*0.229=34.35 150-34.35=115.65 150.00
not have 9850-9830.3=19.7 9850*0.998=9830.30   9850.00
total 54.05 9945.95 10000.00

positive predictive value =P(true positive)/P(total positiive) =34.35/54.05 =0.6355 ~ 63.55 %

negative predictive value =P(true negative)/P(total negative) =9830.3/9945.95=0.9884 ~98.84 %